OpenStudy (loser66):

I have to prove a linear transformation: V is the set C of complex numbers regarded as a real vector space; Ax is the complex conjugate of x But I am not sure about "Complex numbers regarded as a Real vector space" Someone explains me, please

OpenStudy (loser66):

@mathmale

OpenStudy (anonymous):

i think it just means that \(\mathbb{C}\) can be considered as vector space over the real numbers \(\mathbb{R}\) and any element of \(\mathbb C\) can be written as \(a+bi\) where \(a,b\in \mathbb R\)

OpenStudy (anonymous):

i wouldn't worry to much about it, the proof is to show that the map \(T:\mathbb C\to \mathbb C\) via \[a+bi\mapsto a-bi\] is a linear map

OpenStudy (anonymous):

or rather in your notation \(A(a+bi)=a-bi\)

OpenStudy (loser66):

Thank you so much @satellite73

OpenStudy (anonymous):

yw